1. Use the ratio an+1/an to show that the given siquence {an} is strictly increasing or strictly decreasing.

{nn^2/(2n)!}n=1+infinity

1
Expert's answer
2021-05-16T23:05:02-0400

a= "\left(n^{\smash{n^2}}\right)""\/(2n)!"

an+1 = "\left((n+1)^{\smash{(n+1)^2}}\right)/(2(n+1))!"


an+1/an = "\left((n+1)^{\smash{(n+1)^2}}\right) / (2(n+1))!) /"("\left(n^{\smash{n^2}}\right)/(2n)!)" =

= "((2n)! *\left((n+1)^{\smash{(n+1)^2}}\right)" ) / ("\left(n^{\smash{n^2}}\right) * (2n+2)!)" =

= "\left((n+1)^{\smash{n^2+2n+1}}\right) / ((2n+1)*(2n+2)*""\left(n^{\smash{n^2}}\right))" =

="\left((n+1)^{\smash{n^2+2n}}\right)" "\/ (2*(2n+1)*" "\left(n^{\smash{n^2}}\right))" =

=("\left((n+1)^{\smash{n^2+2n-1}}\right) *(1+1/n)/(2*(2+1/n)*""\left(n^{\smash{n^2}}\right)")


Let's calculate the limit an+1/an

"\lim" (n"\to" +"\infty") ("\left((n+1)^{\smash{n^2+2n-1}}\right)*(1+1/n)/2*(2+1/n)*" "\left(n^{\smash{n^2}}\right)" =

="\lim" (n"\to" +"\infty" ) ("\left((n+1)^{\smash{n^2+2n-1}}\right) /(4*""\left(n^{\smash{n^2}}\right)") = +"\infty"


Thats mean that ratio is increasing


Note: "(2n)! \/ (2n+2)! = 1\/((2n+1)*(2n+2))"